OpenStudy (nickcarey2097):

Please help kinda in a hurry Simplify: (sin Θ − cos Θ) − (sin Θ + cos Θ)^2 −sin^2 Θ −cos^2 Θ 0 −2sin(Θ)cos(Θ) − cos(Θ) + sin(Θ) − 1

TheSmartOne (thesmartone):

Hint (a + b)^2 = a^2 + 2ab + b^2 Use that to expand (sin Θ + cos Θ)^2 and then simply add and subtract like terms, if there are any :)

OpenStudy (mathmale):

first of all, please expand (sin theta + cos theta)^2. Note that exponentiation (such as squaring) has to be done before addition/subtraction.

OpenStudy (nickcarey2097):

so a represents sin and b represents cos right?

OpenStudy (nickcarey2097):

(sin+cos)^2=sin^2+2sincos+cos^2

TheSmartOne (thesmartone):

correct :) Now, do you know the identity sin^2Θ + cos^2Θ = 1 Use that to simplify sin^2Θ+2sinΘcosΘ+cos^2Θ

OpenStudy (nickcarey2097):

Not really familiar with it to much kinda new to this stuff 1+sin (2Θ)?

TheSmartOne (thesmartone):

Leave 2sinΘcosΘ as is. Now simplify (sinΘ - cosΘ) - (2sinΘcosΘ + 1)

TheSmartOne (thesmartone):

Since none of the answer choices have sin(2Θ)

OpenStudy (nickcarey2097):

Can I combine the sin and cos in the first parenthesis with the second ones?

TheSmartOne (thesmartone):

no, they aren't like terms

TheSmartOne (thesmartone):

It's like how you can't simplify x + y + 2xy any more further

OpenStudy (nickcarey2097):

-1-cos(Θ)-sin(2Θ)+sin (Θ)?

OpenStudy (nickcarey2097):

I put it into a calculator and thats what I got

TheSmartOne (thesmartone):

yes, but instead replace sin(2Θ) with 2sinΘcosΘ

OpenStudy (nickcarey2097):

-1-cos(Θ)-sin(2sinΘcosΘ)?

OpenStudy (nickcarey2097):

nvm

OpenStudy (nickcarey2097):

-1-cos(Θ)-2sinΘcosΘ+sin (Θ)

TheSmartOne (thesmartone):

Correct That would be D

OpenStudy (nickcarey2097):

Oh so you just re-arrange the like terms like they were like 2x time 2y is also 2y times 2x?

OpenStudy (nickcarey2097):

Awesome it was correct thank you so much!

TheSmartOne (thesmartone):

Anytime! :)